You're considering investing in a company (subject company) whose stock you have valued (vi dividend growth model) at $35. You have researched comparable companies and based on your analysis have concluded that the Price/ Earnings Ratios (P/E) of those companies in the same industry with similar growth and other characteristics of the subject company approximates 8. The EPS of the subject company is 4. Value the company based upon your recent research on market multiples. How does the value change if the P/E is 10?
Answer #1

If PE is 8 Value should be equal to=4*8=32

If PE is 10 Value should be equal to=4*10=40